OpenStudy (anonymous):

What is the solution set of 6x^2 - 24 = 0? a) {-2, 2} b) {-2} c) {2}

OpenStudy (anonymous):

6x^2 = 24 x^2 = 24/6 x^2 = 4 x = \[\sqrt{4}\] x= 2 ------>(c)